OpenStudy (anonymous):

simplify the following fraction... (2/cos)+ (4/5x-1)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{2}{\cos(x)}+\frac{4}{5x-1}\]

OpenStudy (anonymous):

are you supposed to add the fractions?

OpenStudy (anonymous):

\[(2/\cos \theta) + (4/1-\sin \theta)\]

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

\[\frac{2}{\cos(x)}+\frac{4}{1-\sin(x)}=\] \[=\frac{2(1-\sin(x))+4\cos(x)}{\cos(x)(1-\sin(x)}\] is a start

OpenStudy (anonymous):

dont think there are any fancy trig identities here. just multiply out in the numerator

OpenStudy (anonymous):

i got that far...but do you know how to do the rest?

OpenStudy (anonymous):

numerator is \[2-2\sin(x) +4\cos(x)\]

OpenStudy (anonymous):

nothing factors, nothing cancels. i don't see any fancy identity to use to simplify

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

answer is just what you get

OpenStudy (anonymous):

okay so the denominator is cos x (1-sinx)????